OpenStudy (anonymous):

In a collection of dimes and nickels, there are 36 coins. The total value of the coins is $2.90. How many coins of each type are there? dimes=______ nickels=_____

OpenStudy (anonymous):

x=dimes y=nickels x+y=36 10x+5y=290 2x+y=58 2x+y-2x-2y=58-72 y=14 x=36-y=22

OpenStudy (anonymous):

dimes 22 nickels 14 ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks